OpenStudy (anonymous):

Calculus Problem: Find the derivative of the function f(x) = 4 sin x / 4-4cos x

zepdrix (zepdrix):

Hmmm notice how everything has a 4? :O

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

\[\large 4-4cosx=4(1-1cosx)\] If you factor out the 4 from the bottom like this, you might be able to see a nice cancellation before you differentiate :D

zepdrix (zepdrix):

Then it looks likeeee.. quotient rule. Having trouble setting it up? :O

OpenStudy (anonymous):

i can set it up

zepdrix (zepdrix):

Hmmm ok then what part are you getting stuck on? c:

OpenStudy (anonymous):

well i hadn't tried doing the stuff with the four's like what you just did that might be the problem

zepdrix (zepdrix):

\[\large [\frac{ sinx }{ 1-cosx }]'=\frac{ (sinx)'(1-cosx)-sinx(1-cosx)' }{ (1-cosx)^2 }\]

zepdrix (zepdrix):

Hmm yah, a lot of 4's floating around :( that could be tricky.